### v2

# The adiabatic theorem for non-Hermitian quantum systems with real eigenvalues and the complex geometric phase

Abstract
The adiabatic theorem is one of the most interesting and significant theorems in quantum mechanics. However, the adiabatic theorem can fail for general non-Hermitian quantum systems. In this paper, by utilizing the complex geometric phase, the functional calculus for biorthogonal systems and the Grönwall inequality, we prove rigorously that the adiabatic theorem is still valid for diagonalizable non-Hermitian systems with real eigenvalues. The proof also justifies the definition of a complex Berry phase in non-Hermitian systems.
I Introduction
An interesting and important issue of time-dependent quantum systems is the adiabatic theorem. The theorem states that if a quantum system evolves slowly enough under a time-dependent non-degenerate Hamiltonian and the initial state is in the instantaneous eigenstate of the Hamiltonian, then the final state of the system will remain in the instantaneous eigenstate up to a multiplicative phase factor [1]. In 1950, T. Kato gave an elegant but technically involved proof of the theorem [2]. Since its discovery, the adiabatic theorem has attracted much interests from researchers, and many related topics were discussed [3, 4, 5]. In particular, it is intimately related to the studies of geometric phase and quantum computing, with a wide range of applications [6, 7, 8]. However, it should be mentioned that the original discussions on the adiabatic theorem does not involve the Berry phase, which is now usually considered as an indispensable companion with the adiabatic evolution. In fact, it is more than fifty years later when Berry systematically investigated the geometric phase bearing his name [6]. Later, the concept of geometric phase is also extended to the non-abelian case [9].
The discussions of non-Hermitian quantum systems has a long history and recently there is a growing interests in such systems, for their novel features and potential applications [10, 11, 12]. In particular, the dynamics and topology of non-Hermitian systems, such as the skin effect, has attracted much attentions [13, 14, 15, 16, 17, 18, 19, 20]. The original discussions of the adiabatic theorem is within the framework of Hermitian systems. Recently there are discussions on the adiabatic theorem and geometric phase in non-Hermitian systems [21, 22, 23, 24, 25, 26, 27, 28]. A key difference between the adiabatic evolution of Hermitian and non-Hermitian systems is that the adiabatic theorem may fail for the latter ones.
In this paper, we discuss the adiabatic following problem, with the focus on diagonalizable Hamiltonians with real eigenvalues. By utilizing a complex geometric phase, the Grönwall inequality and generalizing Kato’s functional calculus techniques to biorthogonal systems, we prove rigorously that the adiabatic theorem is still valid for non-Hermitian systems with real eigenvalues. The structure of the paper is organized as follows. In section II, we briefly review the adiabatic theorem. In section III, it is shown that the adiabatic theorem is always valid for diagonalizable Hamiltonians with real eigenvalues. In section IV, we give some discussions, show the insights for the result, and conclude this paper. Section V is the appendix.
II Brief review of adiabatic theorem
Usually, the adiabatic following problem is formulated in a Hermitian system. We briefly review the discussions of the theorem, following the way in [2, 29].
Let $H(s)(0\leq s\leq 1)$ be a collection of Hermitian Hamiltonians. Now by setting $s=\frac{t}{T}$, where $t$ is a new variable and $T$ is a parameter. Apparently, $H(\frac{t}{T})$ changes slower than $H(s)$ if $T\geq 1$. Consider the Schrödinger equation of $H(\frac{t}{T})$,
$i\frac{d}{dt}U_{T}(t)=H(\frac{t}{T})U_{T}(t),$
where $U_{T}(t)$ is the abbreviation of $U_{T}(t,0)$, the evolution operator of $H(\frac{t}{T})$. The subscript letter $T$ implies the fact that $U_{T}(t)$ is generally dependent of the parameter $T$. By rewriting the Schrödinger equation using the variable $s=\frac{t}{T}$, one can obtain
$i\frac{d}{ds}U_{T}(s)=TH(s)U_{T}(s).$ (1)
Note that in Eq. (1), we do not bother to make a difference in notation between $U_{T}(t)$ and $U_{T}(s)$, which usually does not cause confusion. The adiabatic theorem discuss the behaviour of $U_{T}(t)$, or equivalently $U_{T}(s)$, when $T$ tends to infinity. A mathematical formalism of the theorem is
$\lim_{T\rightarrow\infty}U_{T}(s)|j(0)\rangle=P_{j}(s)\lim_{T\rightarrow\infty}U_{T}(s)|j(0)\rangle,$ (2)
where $|j(s)\rangle$ is the $j$-th eigenstate of $H(s)$ and $P_{j}(s)$ is the corresponding orthogonal projection. Briefly speaking, when the initial state is set in $|j(0)\rangle$ and evolves slowly enough, the theorem or equivalently Eq. (2) guarantees that the state follows the $j$-th eigenstate $|j\rangle$ up to a factor. On the other hand, in the degenerate case, the state will remain in the subspace spanned by the eigenvectors of the same eigenvalue, which introduced the non-abelian geometric phase [9]. The proof of the adiabatic theorem utilizes the properties of orthogonal projections and unitary operators can be found in [2, 29]. Thus when discussing the non-Hermitian quantum systems, the proof is not directly applicable.
III The main result
III.1 The generalized adiabatic theorem and its proof
In this section, we prove the following result.
The adiabatic theorem is valid for diagonalizable Hamiltonians with real eigenvalues.
Let $H(s)$ be a Hamiltonian satisfying
$H(s)v_{i}(s)=\lambda_{i}(s)v_{i}(s),$ (3)
where $\lambda_{i}(s)$ and $v_{i}(s)$ are the real eigenvalues and corresponding eigenvectors, respectively. Now let $\xi_{i}(s)$ be the corresponding biorthogonal vectors, i.e. $\xi_{i}^{\dagger}(s)v_{i}(s)=\delta_{ij}.$ Note that $\xi_{i}(s)$ always exist. In fact, denote $V(s)$ the matrix whose column vectors are $v_{i}(s)$, then $\xi_{i}^{\dagger}(s)$ are the row vectors of the inverse of $V(s)$. Since the eigenvalues can be degenerate, $H(s)$ and $I$ can be rewritten as
$H(s)=\sum_{k}\lambda_{k}(s)P_{k}(s),$ (4)
$I=\sum_{k}P_{k}(s),$ (5)
where $\lambda_{k}(s)$ are the different eigenvalues and $P_{k}(s)$ are the projections such that $P_{k}(s)P_{l}(s)=\delta_{kl}P_{k}(s)$. To see the existence of $P_{k}(s)$, note that one can take the eigenvectors of the same eigenvalue together. For example, let the eigenvectors of $\lambda_{k}(s)$ be $v_{k_{j}}(s)$, then $P_{k}(s)=\sum_{k_{j}}v_{k_{j}}(s)\xi^{\dagger}_{k_{j}}(s)$. Moreover, it is obvious that $P_{k}(s)$ is independent of the choice of $v_{k_{j}}(s)$. In fact, if one takes $v^{\prime}_{k_{j}}(s)$ and $\xi^{\prime}_{k_{j}}(s)$, then direct calculations show that $P_{k}(s)=\sum_{k_{j}}v_{k_{j}}(s)\xi^{\dagger}_{k_{j}}(s)=\sum_{k_{j}}v^{\prime}_{k_{j}}(s)(\xi^{\prime})^{\dagger}_{k_{j}}(s)$ since $v_{k_{j}}(s)$, $v^{\prime}_{k_{j}}(s)$ are related by an invertible matrix and $\xi_{k_{j}}(s)$, $\xi^{\prime}_{k_{j}}(s)$ are related by its inverse matrix.
Now for any $j$, one can define a matrix $S(s)$ by
$S(s)=\sum_{k\neq j}\frac{1}{\lambda_{k}(s)-\lambda_{j}(s)}P_{k}(s),$ (6)
which satisfies that following relation
$S(s)P_{j}(s)=P_{j}(s)S(s)=0,$ (7)
$(H(s)-\lambda_{j}(s)I)S(s)=I-P_{j}(s).$ (8)
Now suppose the evolution operator is $U_{T}(s)$, then by Eq. (1)
$i\frac{d}{ds}U_{T}^{-1}(s)=-U_{T}^{-1}(s)TH(s).$ (9)
One can further obtain
$\frac{d}{ds}[e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)]$ (10)
$=$ $\frac{d}{ds}[e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}]U_{T}^{-1}(s)+e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}\frac{d}{ds}[U_{T}^{-1}(s)]$
$=$ $-iT\lambda_{j}(s)e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)$
- $e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\frac{d}{ds}[U_{T}(s)]U_{T}^{-1}(s)$
$=$ $-iT\lambda_{j}(s)e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)$
$+$ $e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)iTH(s)$
$=$ $iTe^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)[H(s)-\lambda_{j}(s)I].$
On the other hand, there exists some operator $K_{j}(s)$ satisfying
$\frac{d}{ds}[K_{j}(s)v_{j_{l}}(s)]=[I-P_{j}(s)]\frac{d}{ds}[K_{j}(s)v_{j_{l}}(s)].$ (11)
To see this, note that one can define $K_{j}(s)$ by
$\langle\xi_{i}(s)|K^{\prime}_{j}(s)|v_{k}(s)\rangle=-\sum_{l}\langle\xi_{i}(s)|K_{j}(s)|v_{l}(s)\rangle\langle\xi_{l}(s)|\dot{v}_{k}(s)\rangle.$
Equivalently, the above equation can be written in the matrix form
$K^{\prime}_{j}(s)=-K_{j}(s)G(s),$
where $G_{lk}(s)=\langle\xi_{l}(s)|\dot{v}_{k}(s)\rangle$. The above equation has solution which are often represented by using the time ordering operator. Moreover, one can verify that the above operator satisfies Eq. (11). Moreover, $K_{j}(s)v_{j_{l}}(s)$ lies in the subspace of $\lambda_{j}(s)$.
By substituting Eq. (8) into Eq. (11), we have
$\frac{d}{ds}[K_{j}(s)v_{j_{l}}(s)]$
$=$ $[I-P_{j}(s)]\frac{d}{ds}[K_{j}(s)v_{j_{l}}(s)]$
$=$ $[H(s)-\lambda_{j}(s)I]S(s)\frac{d}{ds}[K_{j}(s)v_{j_{l}}(s)].$
For simplicity, denote $K_{j}(s)v_{j_{l}}(s)$ by $\tilde{v}_{j_{l}}(s)$. Eq. (III.1) can be rewritten as
$\frac{d}{ds}[\tilde{v}_{j_{l}}(s)]=(H(s)-\lambda_{l}(s)I)S(s)\frac{d}{ds}[\tilde{v}_{j_{l}}(s)].$ (14)
Then we have
$\frac{d}{ds}[e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j_{l}}(s)]$ (15)
$=$ $\frac{d}{ds}[e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)]\tilde{v}_{j_{l}}(s)$
$+$ $e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\frac{d}{ds}[\tilde{v}_{j_{l}}(s)]$
By Eqs. (10) and (7), the first term in Eq. (15)
$\frac{d}{ds}[e^{-iT\int_{0}^{\sigma}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)]\tilde{v}_{j_{l}}(s)$ (16)
$=$ $iTe^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)[H(s)-\lambda_{j}(s)I]\tilde{v}_{j_{l}}(s)$
$=$ $0.$
Thus Eq. (15) reduces to
$\frac{d}{ds}[e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j_{l}}(s)]$ (17)
$=$ $e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\frac{d}{ds}[\tilde{v}_{j_{l}}(s)].$
Now by taking integral in Eq. (17),
$e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j_{l}}(s)-\tilde{v}_{j}(0)$
$=$ $\int_{0}^{s}\frac{d}{d\sigma}[e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)\tilde{v}_{j_{l}}(\sigma)]d\sigma$
$=$ $\int_{0}^{s}e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j_{l}}(\sigma)]d\sigma.$
By Eqs. (14), (10) and integration by parts, it follows that
$e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j_{l}}(s)-\tilde{v}_{j}(0)$
$=$ $\int_{0}^{s}e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j_{l}}(\sigma)]d\sigma$
$=$ $\int_{0}^{s}e^{-iT\int_{0}^{s}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)(H(\sigma)-\lambda_{j}(\sigma)I)S(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j_{l}}(\sigma)]d\sigma$
$=$ $\frac{1}{iT}\int_{0}^{s}\frac{d}{d\sigma}[e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)]S(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j_{l}}(\sigma)]d\sigma$
$=$ $\frac{1}{iT}[e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)]S(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j_{l}}(\sigma)]|_{0}^{s}$
- $\frac{1}{iT}\int_{0}^{s}e^{-iT\int_{0}^{\sigma}\lambda_{j}(r)dr}U_{T}^{-1}(\sigma)\frac{d}{d\sigma}[S(\sigma)\frac{d}{d\sigma}[\tilde{v}_{j_{l}}(\sigma)]]d\sigma.$
As we shall prove later, $U_{T}^{-1}(s)$ is uniformly bounded with respect to $T$. Then the Eq. (LABEL:adiabatic1) tends to vanish and we have
$e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}^{-1}(s)\tilde{v}_{j_{l}}(s)-\tilde{v}_{j_{l}}(0)\rightarrow 0.$ (19)
Moreover, as we shall prove later, $U_{T}(s)$ is uniformly bounded with respect to $T$. Thus by multiplying $U_{T}(s)$ on both sides of (19), one can obtain
$U_{T}(s)v_{j_{l}}(0)\rightarrow e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}K_{j}(s)v_{j_{l}}(s).$ (20)
This is exactly the adiabatic theorem. When the initial state is the eigenstate and the system evolves slowly enough, the final state remains in the subspace of instantaneous eigenstates. In particular, when the eigenvalues are non-degenerate, one has
$U_{T}(s)v_{j}(0)\rightarrow e^{-iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle dr}v_{j}(s),$
the state remains an eigenstate up to a factor. The factor consists of two parts, one is the dynamic phase and the other is the Berry phase.
III.2 The uniform boundedness of $U_{T}(s)$ and $U_{T}^{-1}(s)$
We still have to prove $U_{T}(s)$ and $U_{T}^{-1}(s)$ are uniformly bounded with respect to $T$.
For $U_{T}^{-1}(s)$, one can take the adjoint of the Eq. (17), then
$\frac{d}{ds}[\tilde{v}^{\dagger}_{j}(s)e^{iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}[U_{T}^{-1}]^{\dagger}(s)]$ (22)
$=$ $\frac{d}{ds}[\tilde{v}^{\dagger}_{j}(s)]e^{iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}[U_{T}^{-1}]^{\dagger}(s).$
Construct a matrix $\tilde{V}(s)$, whose column vectors are $\tilde{v}_{j}(s)$ and denote $E_{T}(s)$ the diagonal matrix whose diagonal entries are $e^{iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}$. It follows from Eq. (22) that
$\frac{d}{ds}[E_{T}(s)\tilde{V}^{\dagger}(s)[U_{T}^{-1}]^{\dagger}(s)]$ (23)
$=$ $E_{T}(s)\frac{d}{ds}[\tilde{V}^{\dagger}(s)][U_{T}^{-1}]^{\dagger}(s).$
Eq. (23) can be rewritten as
$\frac{d}{ds}[E_{T}(s)\tilde{V}^{\dagger}(s)[U_{T}^{-1}]^{\dagger}(s)]$
$=$ $E_{T}(s)\frac{d}{ds}[\tilde{V}^{\dagger}(s)][\tilde{V}^{\dagger}(s)]^{-1}E^{\dagger}_{T}(s)[E_{T}(s)\tilde{V}^{\dagger}(s)[U_{T}^{-1}]^{\dagger}(s)]$
By taking integral,
$E_{T}(s)\tilde{V}^{\dagger}(s)[U_{T}^{-1}]^{\dagger}(s)-\tilde{V}^{\dagger}(0)$
$=$ $\int_{0}^{s}E_{T}(\sigma)\frac{d}{d\sigma}[\tilde{V}^{\dagger}(\sigma)][\tilde{V}^{\dagger}(\sigma)]^{-1}E^{\dagger}_{T}(\sigma)$
$\cdot[E_{T}(\sigma)\tilde{V}^{\dagger}(\sigma)[U_{T}^{-1}]^{\dagger}(\sigma)]d\sigma.$
By using the triangle inequality of norm and Grönwall inequality (see Appendix B), one can see that
$\|E_{T}(s)\tilde{V}^{\dagger}(s)[U_{T}^{-1}]^{\dagger}(s)\|\leqslant\|\tilde{V}^{\dagger}(0)\|$
$+$ $\int_{0}^{s}\|\tilde{V}^{\dagger}(0)\|\|E_{T}(\sigma)\frac{d}{d\sigma}[\tilde{V}^{\dagger}(\sigma)][\tilde{V}^{\dagger}(\sigma)]^{-1}E^{\dagger}_{T}(\sigma)\|$
$\cdot e^{\int_{\sigma}^{s}\|E_{T}(r)\frac{d}{dr}[\tilde{V}^{\dagger}(r)][\tilde{V}^{\dagger}(r)]^{-1}E^{\dagger}_{T}(r)\|dr}d\sigma.$
Note that both $\tilde{V}(s)$ and $\frac{d}{ds}[\tilde{V}^{\dagger}(s)][\tilde{V}^{\dagger}(s)]^{-1}$ are continuous and independent of $T$, thus they are uniformly bounded. $E_{T}(s)$ is also uniformly bounded. It follows from Eq. (LABEL:norm1) that $\|E_{T}(s)\tilde{V}^{\dagger}(s)[U_{T}^{-1}]^{\dagger}(s)\|$ is uniformly bounded. Thus $U^{-1}_{T}(s)$ is uniformly bounded with respect to $T$.
Along similar lines, one can prove that $U_{T}(s)$ is also uniformly bounded. In fact,
$\frac{d}{ds}[\xi^{\dagger}_{j}(s)e^{iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}(s)]$
$=$ $\frac{d}{ds}[\xi^{\dagger}_{j}(s)]e^{iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}(s)$
$+$ $\xi^{\dagger}_{j}(s)\frac{d}{ds}[e^{iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}(s)]$
Note that by Eqs. (4) and (5),
$\xi^{\dagger}_{j}(s)\frac{d}{ds}[e^{iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}(s)]$
$=$ $\xi^{\dagger}_{j}(s)iTe^{iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}(\lambda_{j}(s)I-H(s))U_{T}(s)$
$=$ $0.$
Hence it follows that
$\frac{d}{ds}[\xi^{\dagger}_{j}(s)e^{iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}(s)]$
$=$ $\frac{d}{ds}[\xi^{\dagger}_{j}(s)]e^{iT\int_{0}^{s}\lambda_{j}(\sigma)d\sigma}U_{T}(s).$
Construct a matrix $\Xi(s)$, whose column vectors are $\xi_{j}(s)$. It follows from Eq. (III.2) that
$\frac{d}{ds}[E_{T}(s)\Xi^{\dagger}(s)U_{T}(s)]$
$=$ $E_{T}(s)\frac{d}{ds}[\Xi^{\dagger}(s)]U_{T}(s)$
$=$ $E_{T}(s)\frac{d}{ds}[\Xi^{\dagger}(s)][\Xi^{\dagger}(s)]^{-1}E^{\dagger}_{T}(s)[E_{T}(s)\Xi^{\dagger}(s)U_{T}(s)]$
Applying the Grönwall inequality,
$\|E_{T}(s)\Xi^{\dagger}(s)U_{T}(s)\|\leqslant\|\Xi^{\dagger}(0)\|$
$+$ $\int_{0}^{s}\|\Xi^{\dagger}(0)\|\|E_{T}(\sigma)\frac{d}{d\sigma}[\Xi^{\dagger}(\sigma)][\Xi^{\dagger}(\sigma)]^{-1}E^{\dagger}_{T}(\sigma)\|$
$\cdot e^{\int_{\sigma}^{s}\|E_{T}(r)\frac{d}{dr}[\Xi^{\dagger}(r)][\Xi^{\dagger}(r)]^{-1}E^{\dagger}_{T}(r)\|dr}d\sigma.$
Since $\Xi(s)$ is continuous and independent of $T$, $E_{T}(s)$ is unitary, a similar discussion shows that $U_{T}(s)$ is uniformly bounded with respect to $T$.
IV Discussions and Conclusion
The above discussion shows that the adiabatic theorem is valid for the diagonalizable Hamiltonians with real eigenvalues. That is, the eigenstate will remain in the subspace when evolving slowly enough. When the energy is non-degenerate, it reduces to the Berry phase.
Apparently, the result is somewhat natural. One may wonder why such a result deserves a specified proof. In fact, any Hamiltonian $H(s)$ with non-degenerate real eigenvalues is similar to a Hermitian Hamiltonian $H_{2}(s)$. If $H(s)=V^{-1}H_{2}(s)V$, then the evolution operator of $U_{T}(s)=V^{-1}U^{(2)}_{T}(s)$, where $U^{(2)}_{T}(s)$ is the evolution operator of $TH_{2}(s)$. Moreover, if $x(s)$ is an eigenvector of $H_{2}(s)$, then $V^{-1}x(s)$ is an eigenvector of $H(s)$. Now since $H_{2}(s)$ is Hermitian, the adiabatic theorem is valid for $H_{2}(s)$. Moreover, since $V$ is time independent and $U_{T}(s)=V^{-1}U^{(2)}_{T}(s)$, one can directly verify that the adiabatic theorem is also valid for $H(s)$. However, when $H(s)=V^{-1}(s)H_{2}(s)V(s)$, i.e. the matrix $V(s)$ is dependent of $s$, one has $U_{T}(s)\neq V^{-1}(s)U^{(2)}_{T}(s)$. In this case, one cannot directly obtain the adiabatic theorem for $H(s)$ from $H_{2}(s)$. Hence a proof is necessary.
Unlike the Hermitian case, the eigenvectors of the Hamiltonian $H(s)$ are usually unnormalized in the non-Hermitian case for convenience. That is, the eigenvectors are not uniquely defined since their norms are undetermined. Consequently, the Berry phases are not uniquely determined either. However, this will bring no essential change to the result. To see this, let $v_{j}(s)$ be the eigenvectors of $H(s)$ and $\xi_{j}(s)$ form a biorthogonal system. Assume that there is another choice of eigenvectors $\psi_{j}(s)$ and $\phi_{i}(s)$ form a biorthogonal system. Now since $H(s)$ has nondegenerate eigenvalues, it is obvious that $\psi_{j}(s)=\mu_{j}(s)v_{j}(s)$ and $\phi_{j}(s)=\frac{1}{\overline{\mu}_{j}(s)}\xi_{j}(s)$, where $\mu_{j}(s)$ is some complex number and $\overline{\mu}_{j}(s)$ is its conjugate.
Now we consider the relation of $v_{j}(s)e^{i\int_{0}^{s}i\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}$ and $\psi_{j}(s)e^{i\int_{0}^{s}i\langle\phi_{j}(\sigma)|\dot{\psi}_{j}(\sigma)\rangle d\sigma}$. In fact,
$\psi_{j}(s)e^{i\int_{0}^{s}i\langle\phi_{j}(\sigma)|\dot{\psi}_{j}(\sigma)\rangle d\sigma}$
$=$ $\mu_{j}(s)v_{j}(s)e^{-\int_{0}^{s}\langle\phi_{j}(\sigma)|\dot{\psi}_{j}(\sigma)\rangle d\sigma}$
$=$ $\mu_{j}(s)v_{j}(s)e^{-\int_{0}^{s}\frac{1}{\mu_{j}(s)}\langle\xi_{j}(\sigma)|[\mu_{j}(\sigma)v_{j}(\sigma)]^{\prime}\rangle d\sigma}$
$=$ $\mu_{j}(s)v_{j}(s)e^{-\int_{0}^{s}\frac{\dot{\mu}_{j}(s)}{\mu_{j}(s)}d\sigma-\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}$
$=$ $\mu_{j}(0)v_{j}(s)e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}.$
Now suppose that one has an initial state (unnormalized vector) $q_{j}(0)=c_{j}(0)v_{j}(0)$. The above derivation and direct calculations show that the adiabatically evolved state $q_{j}(s)$ can always be written as $q_{j}(s)=c_{j}(0)v_{j}(s)e^{-\int_{0}^{s}\langle\xi_{j}(\sigma)|\dot{v}_{j}(\sigma)\rangle d\sigma}$ regardless of using $v_{j}$ or $\psi_{j}$ as the reference in calculation. It should also be mentioned that the Berry phase in our discussion is complex, which is similar to [21]. In fact, a complex Berry phase and the biorthogonal states are also used in [30].
In Eqs. (6)-(8), our proof utilizes the spectral decomposition and functional calculus with respect to biorthogonal systems. Such an approach is inspired by Kato’s discussions on Hermitian systems, which can avoid the Riemann-Lebesgue Lemma in the discussion [2]. However, since its utilization of orthogonal projections and unitary matrices, Kato’s discussions do not directly apply to the non-Hermitian case. Moreover, without the Berry phase or its non-abelian counterpart, a treatment of the differential equation of the orthogonal projections and the intertwining operator is necessary in the proof. This long procedure is not needed in this paper. Due to the use of Berry phase or the non-abelian geometric phase, the proof is largely simplified.
There is one place where this paper is more complex than the Hermitian case. One has to show the uniform boundedness of $U_{T}(s)$ and $U_{T}^{-1}(s)$, which are obvious in Hermitian quantum mechanics. To show the uniform boundedness, we use the Grönwall inequality. Given that our discussion are in a more general case, this complication is not unexpected.
In conclusion, we obtained a more general adiabatic theorem for non-Hermitian systems with real eigenvalues. The discussions also partly justifies the definition of a complex Berry phase and its non-abelian generalization.
V Appendix
V.1 The Grönwall inequality
The integral form of the Grönwall inequality is as follows [31, 32]. Let $I=[a,b)$ be an interval. If $\beta$ is non-negative and $u$ satisfies the following inequality,
$u(t)\leqslant\alpha(t)+\int_{a}^{t}\beta(s)u(s)ds,\forall t\in I,$
then
$u(t)\leqslant\alpha(t)+\int_{a}^{t}\alpha(s)\beta(s)e^{\int_{s}^{t}\beta(r)dr}ds.$
Acknowledgement
This work is partially supported by the National Natural Science Foundation of China (12371135). The first author thanks Prof. Chunlan Jiang, Dianmin Tong and Zeqian Chen for useful suggestions.